# On $3$-manifolds admitting co-orientable taut foliations, but none with vanishing Euler class

### v2

@@ 2. Examples from surgeries on knots in the $3$-sphere
Theorem 1.2 is an immediate consequence of the following two lemmas. The first follows from Ozsváth and Szabó’s proof that a manifold with a co-orientable taut foliation is not an $L$-space [22].
@@
Let $M$ be a rational homology $3$-sphere admitting a co-orientable taut foliation $\mathcal{F}$. Then $M$ has a $\mathrm{spin}^{c}$-structure $\mathfrak{s}$ such that $HF_{\mathrm{red}}(M,\mathfrak{s})\neq 0$ and $c_{1}(\mathfrak{s})=e(\mathcal{F})$.
@@
Let $M$ be a rational homology $3$-sphere that admits a co-orientable taut foliation $\mathcal{F}$. The tangent plane field of $\mathcal{F}$ can be perturbed to obtain a tight contact structure $\xi$ whose contact invariant $c(\xi)\neq 0\in HF_{\mathrm{red}}(M,\mathfrak{s})$, with $c_{1}(\mathfrak{s})=e(\mathcal{F})$ [22]. ∎
The second is an application of the mapping cone formula for calculating the Heegaard Floer homology of surgeries on a knot in the $3$-sphere [23, 24]. In conjunction with Lemma 2.1, this implies Theorem 1.2 since for a 3-manifold $M$ for which $H_{1}(M)$ has odd order, the unique spinc-structure satisfying $c_{1}(\mathfrak{s})=0$ is also the unique self-conjugate spinc-structure on $M$.
@@
Let $M$ be obtained by $p/q$-surgery on a knot $K$ in the $3$-sphere, where $q$ is even and |p/q|>2g(K)1|p/q|>2g(K)-1. Then the unique self-conjugate $\mathrm{spin}^{c}$-structure $\mathfrak{s}$ on $M$ satisfies $HF_{\mathrm{red}}(M,\mathfrak{s})=0$.
@@
Without loss of generality we may assume that $p/q>0$. Let $CFK^{\infty}(K)=\bigoplus_{i,j\in\mathbb{Z}}C\{(i,j)\}$ be the knot Floer complex for $K$. Since |p/q|>2g(K)1|p/q|>2g(K)-1, the mapping cone formula for computing $HF^{+}(M)$ in terms of $CFK^{\infty}(K)$ implies that there is a labelling
@@
such that [21, Proposition 3.5]
$HF_{\mathrm{red}}(M,\phi(i))\cong\bigoplus_{k\in\mathbb{Z}}\bm{A}_{{\rm red},\,\lfloor\frac{pk+i}{q}\rfloor}(K),$
where $\bm{A}_{{\rm red},\,k}(K)$ denotes the reduced part of
$\bm{A}^{+}_{k}=H_{*}(C\{i\geq 0\text{ or }j\geq k\}).$
However, $\bm{A}_{{\rm red},\,k}(K)=0$ for |k|>g(K)1|k|>g(K)-1 [21, Lemma 3.2]. Thus, for $0\leq i\leq p-1$, this becomes
$HF_{\mathrm{red}}(M,\phi(i))\cong\begin{cases}\bm{A}_{{\rm red},\,\lfloor\frac{i}{q}\rfloor}(K),&0\leq i<qg(K)-q,\\ \bm{A}_{{\rm red},\,\lfloor\frac{i-p}{q}\rfloor}(K),&p-qg(K)<i<p,\\ 0,&qg(K)-q\leq i\leq p-qg(K).\end{cases}$
In particular, for each $|j|\leq g(K)-1$, there are precisely $q$ $\mathrm{spin}^{c}$-structures for which $HF_{\mathrm{red}}(M,\mathfrak{s})$ is isomorphic to $\bm{A}_{{\rm red},\,j}(K)$, and for all other $\mathrm{spin}^{c}$-structures, $HF_{\mathrm{red}}(M,\mathfrak{s})=0$. Thus the set
$S=\{\mathfrak{s}\in\mathrm{Spin}^{c}(M)\mid HF_{\mathrm{red}}(M,\mathfrak{s})\neq 0\}$
has cardinality divisible by $q$. In particular, $|S|$ is even. Since conjugation of $\mathrm{spin}^{c}$-structures induces an isomorphism of Heegaard Floer homology, conjugation acts as an involution on $S$. This implies that there are an even number of self-conjugate $\mathrm{spin}^{c}$-structures in $S$. Since $p$ is odd, there is precisely one self-conjugate spinc-structure on $M$, so it follows that $S$ does not contain the self-conjugate spinc-structure on MM. [^2] An alternative line of argument is to use the fact that conjugation of spinc-structures corresponds to the involution $i\mapsto p+q-1-i$ (mod $p$) on $\mathbb{Z}/p$. For $q$ even, this implies that the self-conjugate spinc-structure is labelled by $s=\frac{p+q-1}{2}$. However since |p/q|>2g(K)1|p/q|>2g(K)-1, the above calculation shows that $HF_{\mathrm{red}}(M,\phi(s))=0$.
@@
Let $M$ be a rational homology sphere obtained by integer surgery on a genus one knot in the 3-sphere. Then every co-orientable taut foliation on $M$ has Euler class zero.
@@
Let $K$ be a genus one knot and let $M$ be obtained by $p$-surgery on $K$. Without loss of generality we may assume that $p>0$. The proposition is deduced from Lemma 2.1 by showing that there is at most one spinc-structure $\mathfrak{s}_{0}$ for which $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})\neq 0$ and then verifying that $c_{1}(\mathfrak{s}_{0})=0$.
First we compute $HF_{\mathrm{red}}(M)$. For integer surgeries of slope $p\geq 2g(K)-1$, the mapping cone formula gives a labelling
(2.0.1) $\phi:\mathbb{Z}/p\to\mathrm{Spin}^{c}(M),$
such that [21, Proposition 3.5]
@@
$HF_{\mathrm{red}}(M,\phi(i))\cong\bigoplus_{k\in\mathbb{Z}}\bm{A}_{{\rm red},\,i+kp}(K).$
However, as $g(K)=1$, we have $\bm{A}_{{\rm red},\,k}(K)=0$ for $|k|>0$ [21, Lemma 3.2]. Thus, for $0\leq i\leq p-1$, this becomes
$HF_{\mathrm{red}}(M,\phi(i))\cong\begin{cases}\bm{A}_{{\rm red},\,0}(K),&i=0,\\ 0,&0<i\leq p-1,\end{cases}$
giving that that there is at most one spinc-structure $\mathfrak{s}_{0}$ for which $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})\neq 0$ and that this is the spinc-structure corresponding to 0 under (2.0.1). We verify that $c_{1}(\mathfrak{s}_{0})=0$ by unpacking the labelling in (2.0.1). As in [23, Section 2.4] the map $\phi$ is obtained as follows. Take $X$ to be the trace of $p$-surgery on $K$. Let $\Sigma]\in H_{2}(X)$ denote that the homology class obtained by capping off the core of the 2-handle with a Seifert surface for $K$. If $\mathfrak{s}\in\mathrm{Spin}^{c}(M)$ is the restriction of $\mathfrak{t}\in\mathrm{Spin}^{c}(X)$, then $\phi(i)=\mathfrak{s}$, where $i$ satisfies
$\langle c_{1}(\mathfrak{t}),[\Sigma]\rangle+p\equiv 2i\bmod 2p.$
Thus the spinc-structure with $\mathfrak{s}_{0}=\phi(0)$ is the restriction of any spinc-structure $\mathfrak{t}_{0}$ on $X$ such that $\langle c_{1}(\mathfrak{t}_{0}),[\Sigma]\rangle\equiv p\bmod 2p$. Since the restriction map $H^{2}(X)\rightarrow H^{2}(M)$ is reduction mod $p$, this implies that $c_{1}(\mathfrak{s}_{0})=0$, as required. ∎
@@
As mentioned in the introduction, the line of reasoning in the proof of Proposition 2.3 can easily be extended to show that if $K$ is an almost $L$-space knot and $M$ is obtained by $p$-surgery on $K$ for some $p\geq 2g(K)-1$, then any co-orientable taut foliation on $M$ has Euler class zero.

### v3

@@ 2. Examples from surgeries on knots in the $3$-sphere
In this section we prove our results on surgeries on knots in the $3$-sphere.
2.1. $\mathrm{Spin}^{c}$ structures and co-orientable taut foliations
The set $\mathrm{Spin}^{c}(M)$ of $\mathrm{spin}^{c}$ structures $\mathfrak{s}$ on a rational homology $3$-sphere $M$ corresponds to certain classes of pairs $(\xi,\mathfrak{s}_{0})$, where $\xi$ is an oriented $2$-plane sub-bundle of the tangent bundle of $M$ and $\mathfrak{s}_{0}$ is a spin structure on $M$ [15, proof of Proposition 1]. There is a faithful, transitive action of $H^{2}(M)$ on $\mathrm{Spin}^{c}(M)$ [15, Proposition 1]. The Chern class of a $\mathrm{spin}^{c}$ structure $\mathfrak{s}$, denoted $c_{1}(\mathfrak{s})$, is the Euler class of its oriented plane bundle.
The set $\mathrm{Spin}(M)$ of $\mathrm{spin}$-structures on $M$ is identified with $\{\xi_{0}\}\times\mathrm{Spin}(M)\subset\mathrm{Spin}^{c}(M)$, where $\xi_{0}$ is a trivial bundle. This is precisely the set of $\mathrm{spin}^{c}$ structures $\mathfrak{s}$ satisfying $c_{1}(\mathfrak{s})=0$ [15, Remark (c) page 49]. Reversing the orientation of $\xi$ determines an involution on $\mathrm{Spin}^{c}(M)$, called conjugation, whose fixed point set is $\mathrm{Spin}(M)$ [15, Remark (c) page 49]. Thus there are precisely $|\mathrm{Spin}(M)|=|H_{1}(M;\mathbb{Z}/2)|$ self-conjugate spinc-structures on $M$. We record these remarks in a lemma.
@@ 2.1. $\mathrm{Spin}^{c}$ structures and co-orientable taut foliations
Let $M$ be a rational homology $3$-sphere. Then there are $|H_{1}(M)|$ spinc-structures on $M$ and $|H_{1}(M;\mathbb{Z}/2)|$ self-conjugate spinc-structures. Further, a spinc-structure $\mathfrak{s}$ is self-conjugate if and only if $c_{1}(\mathfrak{s})=0$.
In what follows, we take $\mathbb{F}$ to be $\mathbb{Z}/2$.
Lemma 2.2.
Let $M$ be a rational homology $3$-sphere admitting a co-oriented taut foliation $\mathcal{F}$. Then $M$ has a $\mathrm{spin}^{c}$ structure $\mathfrak{s}$ such that $HF_{\mathrm{red}}(M,\mathfrak{s})$ contains a direct $\mathbb{F}$ summand and $c_{1}(\mathfrak{s})=e(\mathcal{F})$.
@@
Let $M$ be a rational homology $3$-sphere that admits a co-oriented taut foliation $\mathcal{F}$. The tangent plane field of $\mathcal{F}$ can be perturbed to obtain a tight contact structure $\xi$ whose contact invariant $c(\xi)\neq 0\in HF_{\mathrm{red}}(M,\mathfrak{s})$, with $c_{1}(\mathfrak{s})=e(\mathcal{F})$ [36]. Moreover, Lin deduces that the $\mathbb{F}[U]$-module $HF_{\mathrm{red}}(M,\mathfrak{s})$ contains direct $\mathbb{F}$ summand [27]. ∎
Combined, these two lemmas yield the following obstruction to the existence of taut foliations of Euler class zero and non-zero.
Proposition 2.3.
Let $M$ be a rational homology sphere with a co-oriented taut foliation $\mathcal{F}$.
  1. (1)
    If $e(\mathcal{F})=0$, then there is a self-conjugate $\mathrm{spin}^{c}$ structure $\mathfrak{s}$ such that $HF_{\mathrm{red}}(M,\mathfrak{s})$ contains a direct $\mathbb{F}$ summand.
  2. (2)
    If $e(\mathcal{F})\neq 0$, then there is $\mathrm{spin}^{c}$ structure $\mathfrak{s}$ such that $HF_{\mathrm{red}}(M,\mathfrak{s})$ contains a direct $\mathbb{F}$ summand and $\mathfrak{s}$ is not self-conjugate.
In particular, Heegaard Floer homology can be used to show that certain 3-manifolds, including almost $L$-spaces, do not admit co-orientable taut foliations with non-zero Euler class.
Proposition 2.4.
Let $M$ be a rational homology sphere for which there is a unique $\mathrm{spin}^{c}$ structure $\mathfrak{s}_{0}$ for which $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})$ is non-zero. Then any co-orientable taut foliation on $M$ has Euler class zero.
Proof.
Since conjugation of spinc-structures induces an isomorphism on Heegaard Floer homology, we see that $\mathfrak{s}_{0}$ is necessarily self-conjugate. Thus Proposition 2.3 implies that $M$ cannot admit a co-orientable taut foliation with non-zero Euler class. ∎
2.2. Surgery on knots
In order to apply Proposition 2.3, we need examples of manifolds for which we can understand the Heegaard Floer homology. For surgery on knots $K$ in the $3$-sphere, the Heegaard Floer homology can be calculated using the mapping cone formula [38, 39].
Let $CFK^{\infty}(K)=\bigoplus_{i,j\in\mathbb{Z}}C\{(i,j)\}$ be the knot Floer complex of $K$. We will use $\bm{A}_{{\rm red},\,k}(K)$ to denotes the reduced part of
$\bm{A}^{+}_{k}=H_{*}(C\{i\geq 0\text{ or }j\geq k\}).$
The mapping cone formula expresses the reduced Heegaard Floer homology of surgeries on $K$ in terms of the $\bm{A}_{{\rm red},\,k}(K)$ [38, 39]. For our purposes, the following calculation will be sufficient.
@@ 2.2. Surgery on knots
Let $M$ be obtained by $p/q$-surgery on a knot $K$ of genus $g$ in $S^{3}$ with $p/q\geq 2g-1$. Then there is an affine map
$\phi:\mathbb{Z}/p\to\mathrm{Spin}^{c}(M),$
such that for $i$ in the range $\frac{q-p}{2}\leq i\leq\frac{p+q-1}{2}$
$HF_{\mathrm{red}}(M,\phi(i))\cong\begin{cases}\bm{A}_{{\rm red},\,\left\lfloor\frac{i}{q}\right\rfloor}(K),&q(1-g)\leq i\leq qg-1,\\ 0,&\text{otherwise}.\end{cases}$
@@
Since $p/q\geq 2g-1$, the mapping cone formula for computing $HF^{+}(M)$ in terms of $CFK^{\infty}(K)$ implies that there is a labelling
@@
such that [35, Proposition 3.5]
$HF_{\mathrm{red}}(M,\phi(i))\cong\bigoplus_{k\in\mathbb{Z}}\bm{A}_{{\rm red},\,\lfloor\frac{pk+i}{q}\rfloor}(K).$
Each value of $i$ (mod $p$) is represented for $i$ in the range $\frac{q-p}{2}\leq i\leq\frac{p+q-1}{2}$, so as $\bm{A}_{{\rm red},\,j}(K)=0$ for $|j|\geq g$ [35, Lemma 3.2], we have
$HF_{\mathrm{red}}(M,\phi(i))\cong\begin{cases}\bm{A}_{{\rm red},\,\lfloor\frac{i}{q}\rfloor}(K),&q(1-g)\leq i\leq qg-1,\\ 0&\text{otherwise}.\end{cases}$
Combined, Proposition 2.3 and Lemma 2.5 quickly yield Theorem 1.3.
Proof of Theorem 1.3.
Since $p$ is odd, there is an odd number of spinc-structures on $M$ of which exactly one, call it $\mathfrak{s}_{0}$, is self-conjugate (Lemma 2.1). After possibly reversing orientation on $M$ and replacing $K$ by its mirror image, we may assume that $p/q\geq 2g(K)-1$. By Lemma 2.5, we see that for each $j$ such that $|j|<g(K)$, there are precisely $q$ spinc-structures for which $HF_{\mathrm{red}}(M,\mathfrak{s})\cong\bm{A}_{{\rm red},\,j}(K)$. Thus the number of spinc-structures for which $HF_{\mathrm{red}}(M,\mathfrak{s})\neq 0$ is divisible by $q$ and hence even. Consequently, we see that there are an odd number of spinc-structures such that $HF_{\mathrm{red}}(M,\mathfrak{s})=0$. Since conjugation of $\mathrm{spin}^{c}$ structures induces an isomorphism of Heegaard Floer homology, this implies that $\mathfrak{s}_{0}$ on $M$ satisfies $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})=0$. By Proposition 2.3, this implies that no co-orientable taut foliation on $M$ can have Euler class zero. ∎
@@
Let $M$ be a rational homology sphere obtained by $p$-surgery on a knot $K$ is the $3$-sphere, for some $p\in\mathbb{Z}$. If $K$ has genus one, or $K$ is an almost $L$-space knot and $p\geq 2g(K)-1$, then any co-orientable taut foliation on $M$ has Euler class zero.
@@
First suppose that $K$ is an almost $L$-space knot. Then $K(p)$ is an almost $L$-space for each integer $p\geq 2g(K)-1$ [8] and as such, there is a unique $\mathrm{spin}^{c}$ structure $\mathfrak{s}_{0}$ for which $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})\neq 0$. Hence, Proposition 2.4 implies that every co-orientable taut foliation on $M$ has Euler class zero.
Next suppose that $K$ is a genus one knot and let $M$ be obtained by $p$-surgery on $K$. Without loss of generality we may assume that $p>0$. In particular, we have $p\geq 2g(K)-1=1$ and we may apply Lemma 2.5 to calculate $HF_{\mathrm{red}}(M)$. This shows that $HF_{\mathrm{red}}(M,\phi(i))\cong 0$ for all $i\not\equiv 0\;(\bmod\;p)$. Thus there is at most one spinc-structure $\mathfrak{s}_{0}$ with $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})\neq 0$. By Proposition 2.4, this implies that every co-orientable taut foliation on $M$ has Euler class zero. ∎
2.3. Negative surgeries on $L$-space knots
Let $K$ be a knot in the 3-sphere. Its symmetrized Alexander polynomial can be written in the form
$\Delta_{K}(t)=\sum_{k}a_{k}t^{k},$
where $a_{k}=a_{-k}$ for all $k$. For $i\geq 0$ we will use $t_{i}(K)$ to denote $i$th torsion coefficient
$t_{i}(K)=\sum_{k\geq 1}ka_{i+k}.$
In what follows we will use $\mathcal{T}(n)$ to denote the cyclic $\mathbb{F}[U]$-module
$\mathcal{T}(n)=\frac{\mathbb{F}[U]}{(U^{n})}.$
Lemma 2.7.
Let $M$ be obtained by p/q-p/q-surgery on a non-trivial $L$-space knot $K$ for $p/q>0$. If $q$ is odd, then $M$ has a self-conjugate spinc-structure $\mathfrak{s}_{0}$ such that
$HF_{\mathrm{red}}(M,\mathfrak{s}_{0})\cong\mathcal{T}(t_{0})\oplus\bigoplus_{k\geq 1}\mathcal{T}(t_{\lfloor\frac{2pk+q-1}{2q}\rfloor})\oplus\mathcal{T}(t_{\lfloor\frac{2pk+q-1}{2q}\rfloor}).$
If $p+q$ is odd, then $M$ has a self-conjugate spinc-structure $\mathfrak{s}_{1}$ such that
$HF_{\mathrm{red}}(M,\mathfrak{s}_{1})\cong\bigoplus_{k\geq 0}\mathcal{T}(t_{\lfloor\frac{2pk+p+q-1}{2q}\rfloor})\oplus\mathcal{T}(t_{\lfloor\frac{2pk+p+q-1}{2q}\rfloor})$
Proof.
When applied to negative surgeries on $L$-space knots [13, Lemma 18], the mapping cone formula implies that there is a labelling of spinc-structures
$\phi:\mathbb{Z}/p\rightarrow\mathrm{Spin}^{c}(M)$
such that
@@ 2.3. Negative surgeries on $L$-space knots
(2.3.1) $HF_{\mathrm{red}}(M,\phi(i))\cong\bigoplus_{k\in\mathbb{Z}}\mathcal{T}(t_{|\lfloor\frac{pk+i}{q}\rfloor|}),$
where we are using that the fact that for $L$-space knots, the quantities $H_{k}$ and $V_{k}$ are determined by the torsion coefficients of the Alexander polynomial:
$V_{k}=H_{-k}=t_{k}(K).$
Using the identity
$\left\lfloor\frac{pk+i}{q}\right\rfloor=-\left\lfloor\frac{q-1-i-pk}{q}\right\rfloor$
alongside (2.3.1) shows that $HF_{\mathrm{red}}(M,\phi(i))\cong HF_{\mathrm{red}}(M,\phi(\tau(i)))$, where $\tau$ is the involution $\tau(i)=q-1-i\;(\bmod\;p)$ on $\mathbb{Z}/p$. The fixed points of $\tau$ are
$\left\{\frac{q-1}{2},\frac{p+q-1}{2}\right\}\cap\mathbb{Z}.$
If $q$ is odd, then we take $\mathfrak{s}_{0}=\phi(\frac{q-1}{2})$. If $p+q$ is odd, then we take $\mathfrak{s}_{1}=\phi(\frac{p+q-1}{2})$. Note that $\mathfrak{s}_{0}$ and $\mathfrak{s}_{1}$ are both defined if and only if $p$ is even and that precisely one of them is defined if $p$ is odd. For these spinc-structures (2.3.1) reduces to
$HF_{\mathrm{red}}(M,\mathfrak{s}_{0})\cong\mathcal{T}(t_{0})\oplus\bigoplus_{k\geq 1}\mathcal{T}(t_{\lfloor\frac{2pk+q-1}{2q}\rfloor})\oplus\mathcal{T}(t_{\lfloor\frac{2pk+q-1}{2q}\rfloor})$
and
$HF_{\mathrm{red}}(M,\mathfrak{s}_{1})\cong\bigoplus_{k\geq 0}\mathcal{T}(t_{\lfloor\frac{2pk+p+q-1}{2q}\rfloor})\oplus\mathcal{T}(t_{\lfloor\frac{2pk+p+q-1}{2q}\rfloor}).$
Since $K$ is non-trivial, $t_{0}(K)\geq 1$ which implies that $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})$ contains an odd number of cyclic summands whereas $HF_{\mathrm{red}}(M,\mathfrak{s}_{1})$ contains an even number of cyclic summands. So $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})\not\cong HF_{\mathrm{red}}(M,\mathfrak{s}_{1})$.
Suppose that $\mathfrak{s}_{i}$ is defined for $i=0$ or $1$. It follows from the action of $\tau$ that the set $S_{i}$ of spinc-structures
$S_{i}=\{\mathfrak{s}\in\mathrm{Spin}^{c}(M)\mid HF_{\mathrm{red}}(M,\mathfrak{s})\cong HF_{\mathrm{red}}(M,\mathfrak{s}_{i})\}$
has odd cardinality. Since Heegaard Floer is invariant under conjugation, it follows that $S_{i}$ contains a self-conjugate spinc-structure. Since $M$ has one self-conjugate spinc-structure if $p$ is odd and two if $p$ is even, it follows that the reduced Heegaard Floer groups in the self-conjugate spinc-structures are isomorphic to $HF_{\mathrm{red}}(M,\mathfrak{s}_{0})$ and $HF_{\mathrm{red}}(M,\mathfrak{s}_{1})$, according to their existence. ∎
For $K$ a non-trivial $L$-space knot of genus $g(K)$, the non-zero coefficients of $\Delta_{K}(t)$ are equal to $\pm 1$ and alternate in sign. That is the symmetrized Alexander polynomial can be written in the form
$\Delta_{K}(t)=\sum_{k=-\ell}^{\ell}(-1)^{k+\ell}t^{\beta_{k}}$
for some sequence of integers $\beta_{l}>\dots>\beta_{-l}$ such that $\beta_{k}=-\beta_{-k}$, $\beta_{\ell}=g(K)$ and $\beta_{\ell-1}=g(K)-1$. From such a description, one sees that the torsion coefficients form a non-increasing sequence such that $t_{g}(K)=0$ and $t_{g-1}=1$. It follows that if one takes $T_{1}$ to be the quantity
$T_{1}=|\{i\mid t_{i}(K)=1\}|,$
then $t_{i}(K)=1$ if and only if
$g-T_{1}\leq i\leq g-1.$
Theorem 2.8.
Let $M$ be obtained by p/q-p/q-surgery on an $L$-space knot $K$ for $p/q>0$. Then $M$ admits a co-orientable taut foliation. However, if $M$ admits a co-orientable taut foliation with Euler class zero, then
q(2g(K)1)=r+mpq(2g(K)-1)=r+mp
for some integers $m\geq 0$ and $1\leq r<2qT_{1}$, where $T_{1}=|\{i\mid t_{i}(K)=1\}|$.
Proof.
An $L$-space knot is fibred and its monodromy has positive fractional Dehn twist coefficient [34, 16]. Consequently Roberts [40] shows that any negative surgery on an $L$-space knot admits a co-orientable taut foliation.
Now suppose that $M$ admits a co-orientable taut foliation with Euler class zero. From Lemma 2.7, Proposition 2.3 and the discussion prior to this proposition, we see that there is a non-negative integer $m$ such that pm+q1pm+q-1 is even and
$g-T_{1}\leq\left\lfloor\frac{pm+q-1}{2q}\right\rfloor\leq g-1.$
Since pm+q1pm+q-1 is even, $\frac{pm+q-1}{2q}$ can be expressed as a fraction with denominator $q$ and so we have
$g-T_{1}\leq\frac{pm+q-1}{2q}\leq g-\frac{1}{q}.$
Rearranging this implies that
$q(2g-1)-2qT_{1}+1\leq pm\leq q(2g-1)-1.$
This implies that there is $1\leq r\leq 2qT_{1}-1$ such that
q(2g1)=r+mp.q(2g-1)=r+mp.
2.4. Applications of Theorem 2.8
Using Theorem 2.8, we can generate many examples of manifolds admitting a co-orientable taut foliation, but none with Euler class zero. For instance, it allows us to drop the condition that $q$ be even from Theorem 1.3 for many $L$-space knots.
Corollary 2.9.
Let $M$ be a manifold obtained by p/q-p/q-surgery on an $L$-space knot $K$ with $t_{0}(K)\geq 2$, where $p/q\geq 2g(K)-1$. Then $M$ admits a co-orientable taut foliation but none with Euler class zero.
Proof.
If $M$ admits a co-orientable taut foliation with zero Euler class then Theorem 2.8 implies that q(2g(K)1)=r+mpq(2g(K)-1)=r+mp for integers $m\geq 0$ and $r$, where $1\leq r<2qT_{1}$. If $m>0$, then p/q=(2g(K)1)/mr/qm<2g(K)1p/q=(2g(K)-1)/m-r/qm<2g(K)-1, contrary to our hypotheses. Thus $m=0$ and therefore $q(2g(K)-1)=r<2qT_{1}$. On the other hand, the condition that $t_{0}(K)\geq 2$ implies that $T_{1}\leq g(K)-1$, so q(2g(K)1)=r<2q(g(K)1)q(2g(K)-1)=r<2q(g(K)-1), which is impossible. Thus $M$ admits no co-orientable taut foliation with Euler class zero. ∎
@@
The condition $t_{0}(K)\geq 2$ appears to be generic among $L$-space knots. The authors are aware of only five $L$-space knots for which it does not hold: the three torus knots $T(2,3)$, $T(2,5)$ and $T(3,4)$, the $(2,3)$-cable of $T(2,3)$, and the $(2,5)$-cable of $T(2,3)$, of genus $1,2,3,3,$ and $4$, respectively. See [26]. Since $t_{0}(K)=1$ if and only if $\Delta_{K}(t)=(t^{g}+t^{-g})-(t^{g-1}+t^{-(g-1)})+1$, where $g=g(K)$, it follows from [42, Corollary 4.5] that if $K$ has a lens space surgery (so, conjecturally, is a Berge knot) and $g(K)\geq 4$ then $t_{0}(K)\geq 2$.
It will be convenient to note that one can bound the quantity $T_{1}$ appearing in Theorem 2.8 using the bridge index.
Corollary 2.11.
Let $K$ be an $L$-space knot of bridge index $\mathrm{br}(K)$ and let $M$ be obtained by p/q-p/q-surgery on $K$ for $p/q>0$. Then $M$ admits a co-orientable taut foliation. However, if $M$ admits a co-orientable taut foliation with Euler class zero, then
q(2g(K)1)=r+mpq(2g(K)-1)=r+mp
for some integers $m\geq 0$ and $1\leq r<2q\mathrm{br}(K)$.
Proof.
Given Theorem 2.8, it suffices to prove that $T_{1}\leq\mathrm{br}(K)$. This follows from the properties of the invariant $\mathrm{Ord}_{v}(K)$ defined in [20]. The symmetrized Alexander polynomial of a non-trivial $L$-space knot can be written in the form
$\Delta_{K}(t)=t^{g}-t^{g-1}+t^{g-T_{1}}+\text{lower order terms}$
The calculation of $\mathrm{Ord}_{v}(K)$ for $L$-space knots [20, Lemma 5.1] shows that
$T_{1}\leq\mathrm{Ord}_{v}(K)+1.$
However, for any knot $K$ the bound $\mathrm{Ord}_{v}(K)+1\leq\mathrm{br}(K)$ holds [20, Corollary 1.9]. ∎
Theorem 1.7.
For any integer $p\geq 7$ there are infinitely many hyperbolic 3-manifolds $M$ such that $H^{2}(M)\cong\mathbb{Z}/p$ and $M$ admits a co-orientable taut foliation, but none with Euler class zero. Further, $\pi_{1}(M)$ is left-orderable.
Proof.
For $g\geq 5$, let $K_{g}$ denote the (2,3,2g3)(-2,3,2g-3)-pretzel knot. The knot $K_{g}$ is a hyperbolic $L$-space knot [37]. Moreover, we have $br(K_{g})=3$ and $g(K)=g$. Let $M$ be obtained by p/q-p/q-surgery on $K_{g}$ for $p/q>0$. The surgery slope 4g14g-1 is an exceptional surgery slope for $K_{g}$ since it yields a Seifert fibred space [4, Proposition 16]. Since $\Delta(4g-1,-p/q)=|(4g-1)q+p|\geq 20$, the known bounds on exceptional fillings imply that $M$ is a hyperbolic manifold [29]. The work of Roberts shows that there exists a co-orientable taut foliation on $M$. Moreover, if $g\neq 6$, then $M$ has left-orderable fundamental group [45, 46].
On the other hand, if $M$ admits a co-orientable taut foliation with Euler class zero, then Corollary 2.11 implies
$q(2g(K_{g})-1)\equiv r\;(\bmod\;p)$
where $1\leq r\leq 6q-1$. Thus we see that $M$ does not admit a co-orientable taut foliation with Euler class zero, whenever $p/q\geq 7$ and $g\equiv 0\;(\bmod\;p)$. ∎